One - Dimensional Kinematics


1. The table below is a table of data from an experiment measuring the variation of velocity with time for a car.
Time (s). 0 10 20 30 40 50

Velocity (m/s)15 20 25 30 35 39
Graph the data on a velocity-time graph.

2. A ripe mango, hanging from a branch, suddenly falls down to the ground. If the mango acquired a final velocity of 8 m/s when it hit the ground?

3. If a ball is released to fall down from the top of a building 30 meters high, how many seconds will it take for the ball to reach the ground?
4. Find the time it takes to travel 30 km for a bus which accelerates uniformly from 40 km/hr to 80 km/hr.
5. A car initially running at 5 m/s uniformly speeds up to 9 m/s over 20 seconds. What is its
acceleration?
1
Expert's answer
2021-09-28T11:14:16-0400

1. The graph is below:



2. The time was


"t=\frac vg=0.82\text{ s}."

3. The time is


"t=\sqrt{2h/g}=2.47\text{ s
}."

4. The time is


"t=\frac{2d}{v_1+v_2}=1\text{ h}."

5. The acceleration is


"a=\frac{v_2-v_1}{t}=0.2\text{ m/s}^2."